Tom is considering buying a $20,000 car. After five years, he will be able to sell the vehicle for $10,000. Gasoline costs will be $2000 per year, insurance $600 per year, and parking $400 per year. Maintenance costs for the first year will be $1000, rising by $200 per year thereafter. Assume all costs are payable at the end of the year. The alternative is for Tom to take taxis everywhere. This will cost an estimated $6000 per year. Tom will rent a vehicle each year at a total cost (to year end) of $600 for the family vacation, if he has no car. If Tom values money at 6%/year (compounded annually), should he buy the car?
Answer #1

EUAC of buying the car = 20000 * (A/P, 6%,5) + 2000 + 600 + 400 + 1000 + 200 * (A/G, 6%,5) - 10000 * (A/F, 6%,5)

= 20000 * 0.237396 + 2000 + 600 + 400 + 1000 + 200 * 1.883633 - 10000 * 0.177396

= 7350.69

EUAC of not buying car = 6000 + 600 = 6600

As EUAC of not buying car is less, he should not buy the car